OpenStudy (ajanijones):

A little help please medals to. What are the solutions to the equation 5 = |x|? Choose all answers that are correct. A. -1/5 B. -5 C. 5 D. 1/5

OpenStudy (ajanijones):

I dont understand how to do this mind explaining it to me,

OpenStudy (anonymous):

well the IxI means its positve so it can't be a or b

OpenStudy (ajanijones):

But A and B are Negative :/

OpenStudy (anonymous):

yeah the solution has to be positive

OpenStudy (ajanijones):

Okay but then why are they negative.

OpenStudy (amistre64):

recall that for some positive integer; k |k| = k and |-k| = k

OpenStudy (ajanijones):

Okay so Its gonna be -5?

OpenStudy (amistre64):

an aboslute value function inputs a number and strips off its sign ...

OpenStudy (amistre64):

which options, when you strip them of sign, equate to (become) 5 ?

OpenStudy (anonymous):

the absolute value of a number is always positive...

OpenStudy (ajanijones):

Okay so if it was -5 = |5|?

OpenStudy (amistre64):

|5|, we strip the sign off of 5 and it becomes ... 5

OpenStudy (amistre64):

so no ... |5| is not equal to -5

OpenStudy (ajanijones):

So then the answers are C. and D.

OpenStudy (anonymous):

yes.

OpenStudy (amistre64):

1/5 does not strip a sign and become 5 ...

OpenStudy (ajanijones):

Okay so now I think I get it so if its 5 = |x| then it will be positive

OpenStudy (amistre64):

thats incorrect

OpenStudy (freckles):

Would x^2=25 be easier for you to solve?

OpenStudy (amistre64):

apparently its not going to be eaiser

OpenStudy (amistre64):

what does an absolute value function do?

OpenStudy (ajanijones):

, An absolute value function is a function that contains an algebraic expression within absolute value symbols. Recall that the absolute value of a number is its distance from 0 on the number line. To graph an absolute value function, choose several values of x and find some ordered pairs.

OpenStudy (amistre64):

ok, and your understanding of that is evidently in error ... so simplify it an absolute value function strips away the sign

OpenStudy (amistre64):

|x| = 5 means that for some signed version of 5, the function stripped it what are the only signs that 5 can have?

OpenStudy (ajanijones):

So it would be C,B because it takes away the Negative.

OpenStudy (amistre64):

good |5| becomes 5 and |-5| becomes 5

OpenStudy (ajanijones):

YAY! :) I like that example you just wrote do you mind if I copy it and save it?

OpenStudy (amistre64):

i got no issue with it :)

OpenStudy (ajanijones):

Okay thank,you to every one.

OpenStudy (freckles):

the reason I brought up the equation equation is because if you square both sides of |x|=5 you get x^2=25 and notice that x=5 or x=-5 are also solutions to x^2=25 since (5)^2=(5)(5)=25 and (-5)^2=(-5)(-5)=25

OpenStudy (freckles):

the other equation*

OpenStudy (amistre64):

shshshsh, its ok ... its ok... everything will be ok :)

OpenStudy (freckles):

just bringing it up just in case he might find it useful later